# Multilevel neural simulation-based inference

### v1

@@ C.5 Cosmological simulations
We use the CAMELS simulation suite (Villaescusa-Navarro et al., 2021, 2023), a benchmark dataset for machine learning in astrophysics, to study multi-fidelity simulation-based inference in cosmology. The dataset includes both low-fidelity (gravity-only N-body) and high-fidelity (hydrodynamic) simulations of 25 Mpc$/h$ cosmological volumes. The original inference task involves two cosmological parameters, $\theta=(\Omega_{m},\sigma_{8})$, where $\Omega_{m}$ is the matter density and $\sigma_{8}$ the amplitude of fluctuations, with mock power spectrum measurements $P(k)$ used to infer the parameters. However, due to the limited availability of both high- and low-fidelity simulations, we focus on inferring only $\sigma_{8}$ and treat $\Omega_{m}$ as part of the nuisance parameter. We found that attempting to jointly infer both parameters led to poor performance (expected due to physical $\sigma_{8}$-$\Omega_{m}$ degeneracy) even when using 90% of the high-fidelity data.
In this setup, low-fidelity simulations are governed solely by $\theta$ and the initial condition seed $u$, while high-fidelity simulations additionally incorporate complex astrophysical processes (e.g., feedback), modelled via four extra parameters. These high-fidelity simulations are significantly more computationally expensive—often more than 100× slower to generate than low-fidelity ones. Given these limitations, cosmological analyses often rely on conservative data cuts to exclude small-scale modes (e.g., $k\gtrsim 0.1,h/\text{Mpc}$), where simulation inaccuracies are most pronounced (e.g. Jeffrey et al., 2025; Gatti et al., 2025). Multi-fidelity approaches aim to mitigate such constraints by leveraging both inexpensive and high-fidelity simulations to improve the accuracy of cosmological inference.
Note that the idea of combining low- and high- fidelity simulations has previously been proposed in cosmology; see for example the work of Chartier et al. (2021); Chartier and Wandelt (2022) who use low-fidelity simulations to construct approximations of quantities of interest which can be used as control variates. This work differs from our proposed approach in that it is based on regression and approximations of quantities such as means and covariances of interest, rather than of the training objective of neural SBI.
@@
Figure 11: Additional results for the cosmological simulator experiment. The one with noise is used to train the neural networks. (a) 1 seed-matched sample from high and low fidelity simulators. (b)-(c) Recovery plot of ML-NPE and NPE. Adding low fidelity samples lead better recovery of the parameter.
Neural network details.
We use NSF with $3$ bins, span of $[-3,3]$ and $3$ coupling layers. The conditioner for the NSF is an MLP with $2$ hidden layer of $30$ units, and $10\%$ dropout, trained for $400$ epochs.
Evaluations details.

### v2

@@ B.5 Cosmological simulations
We use the CAMELS simulation suite [Villaescusa-Navarro et al., 2021, 2023], a benchmark dataset for machine learning in astrophysics, to study multi-fidelity simulation-based inference in cosmology. The simulation is one of the most expensive cosmological suites ever run; a small fraction of the next generation are being run on the UK DIRAC HPC Facility with 15M CPUh.
The dataset includes both low-fidelity (gravity-only N-body) and high-fidelity (hydrodynamic) simulations of 25 Mpc/h/h cosmological volumes. The original inference task involves two cosmological parameters, $\theta=(\Omega_{m},\sigma_{8})$, where $\Omega_{m}$ is the matter density and $\sigma_{8}$ the amplitude of fluctuations, with mock power spectrum measurements $P(k)$ used to infer the parameters. However, due to the limited availability of both high- and low-fidelity simulations, we focus on inferring only $\sigma_{8}$ and treat $\Omega_{m}$ as part of the nuisance parameter. We found that attempting to jointly infer both parameters led to poor performance (expected due to physical $\sigma_{8}$-$\Omega_{m}$ degeneracy) even when using 90% of the high-fidelity data.
In this setup, low-fidelity simulations are governed solely by $\theta$ and the initial condition seed $u$, while high-fidelity simulations additionally incorporate complex astrophysical processes (e.g., feedback), modelled via four extra parameters. These high-fidelity simulations are significantly more computationally expensive—often more than 100× slower to generate than low-fidelity ones. Given these limitations, cosmological analyses often rely on conservative data cuts to exclude small-scale modes (e.g., $k\gtrsim 0.1,h/\text{Mpc}$), where simulation inaccuracies are most pronounced [e.g. Jeffrey et al., 2025, Gatti et al., 2025]. Multi-fidelity approaches aim to mitigate such constraints by leveraging both inexpensive and high-fidelity simulations to improve the accuracy of cosmological inference.
Note that the idea of combining low- and high- fidelity simulations has previously been proposed in cosmology; see for example the work of Chartier et al. [2021], Chartier and Wandelt [2022] who use low-fidelity simulations to construct approximations of quantities of interest which can be used as control variates. This work differs from our proposed approach in that it is based on regression and approximations of quantities such as means and covariances of interest, rather than of the training objective of neural SBI.
Neural network details.
We use NSF with $3$ bins, span of [3,3][-3,3] and $3$ coupling layers. The conditioner for the NSF is an MLP with $2$ hidden layer of $30$ units, and $10\%$ dropout, trained for $400$ epochs.
Evaluations details.
@@
Figure 9: Additional results for the cosmological simulator experiment. (a) 1 seed-matched sample from high and low fidelity simulators. The one with noise is used to train the neural networks. (b)-(c) Recovery plot of ML-NPE and NPE. Adding low fidelity samples lead better recovery of the parameter.
B.6 Ablation study of the gradient adjustment technique
We conduct experiments to evaluate the effect of the different gradient adjustment techniques we employ on the performance of our method. We train both ML-NPE and ML-NLE using (i) our gradient adjustment approach which involves both rescaling and projection, (ii) only rescaling, (iii) only projection, and (iv) no gradient adjustment (standard training). We then compare the performance of NPE and NLE on the g-and-k and NLE on the toggle switch experiment. Other than the choice of the gradient adjustment, the training method, hyperparameters, and evaluation methods remain the same. The results are shown in Table 4.
Table 4: Mean and standard deviation of the metrics. For g-and-k NPE, $n_{0}=1000,n_{1}=100,m=1000$ and for g-and-k NLE, $n_{0}=10^{4},n_{1}=300,m=1$ as in Section˜5.1. For toggle switch NLE, $n_{0}=10000,n_{1}=500,n_{2}=100$ as in Section˜5.3.
(i) both (ii) only rescaling (iii) only projection (iv) standard
g-and-k NPE (NLPD $\downarrow$) -0.30 -0.21 -0.11 -0.13
(0.31) (0.25) (0.11) (0.27)
g-and-k NLE (KLD $\downarrow$) 0.22 0.21 0.24 0.24
(0.47) (0.45) (0.46) (0.28)
Toggle-switch NLE (MMD $\downarrow$) 0.26 0.50 0.35 0.59
(0.25) (0.37) (0.27) (0.31)
We observe that when the MLMC loss diverges under standard training, as is the case for the toggle switch experiment and g-and-k with NPE, our gradient adjustment approach that combines both gradient rescaling and projection yields the best results. In the case of NLE on the g-and-k simulator, the loss does not diverge and standard training is sufficient. In such cases, the gradient adjustment yields similar results as standard training, albeit with an increase in the variance of the metric. Therefore, we suggest optimising using our gradient adjustment approach when using ML-NLE or ML-NPE as it avoids the need to first detect whether the loss is diverging or not.
We further compare the training losses with and without gradient adjustment. With the gradient adjustment, the loss diverges, whereas with it, the loss curve exhibits convergence; see Figure˜10. To clarify the reason for this behaviour, we also provide an illustration of our gradient scaling procedure in Figure˜11.
We note that the optimisation requires some form of regularisation, and the one we used is one possible choice. In our experiments, the gradient adjustment approach performed the best among the options we tried (e.g., regularising the contribution of the difference term when it dominates the first term, or penalising high variance in the difference term). Moreover, it has the added benefit of not requiring any additional hyperparameters.
((a)) Loss without gradient adjustment
((b)) Loss with gradient adjustment
Figure 10: Comparison of training losses with and without gradient adjustment. (a) Without gradient adjustment, the contribution of $\zeta_{\phi}^{0,-}$ begins to dominate the loss around epoch 1000. The resulting conflict between gradient components leads to unstable optimisation, as evidenced by strong fluctuations in the loss and eventual divergence. (b) With gradient adjustment, all components contribute more stably, and the overall loss decreases steadily eventually, indicating convergence. Loss functions are shown for the g-and-k experiment with NLE.
((a))
((b))
((c))
[…]